OpenStudy (anonymous):

solve for x 10/3x + 4/3 = 7+x/2x

OpenStudy (pfenn1):

\[\left( 10 \over 3x \right) +\left( 4 \over 3 \right) = \left( 7 +x \over 2x \right)\]

OpenStudy (pfenn1):

Is that the correct equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

are you solving it?

OpenStudy (anonymous):

no solve it by your own

OpenStudy (anonymous):

HOW!?!?!

OpenStudy (pfenn1):

Find the common denominator

OpenStudy (anonymous):

now in this how u want to get the solution right

OpenStudy (anonymous):

first multiply both sides by 6 and tell me what u are getting